if the minimum horizontal  force required to drag a heavy body on a rough horizontal surface be n times its magnitude, when it be inclined to the horinzontal at the angle of friction. find coefficient of friction.


 


 


  

Grade:12

1 Answers

Ramesh V
14 years ago

The free body diagram is shown below:

on balancing forces we have

F and R are forces and u (=tan x) is coeff of friction

N1 =mg  and F = f1 = uN1 =umg

N2 = mg.cosx

f2=uN2=umgcos x

R=f2+mgsinx

R = mg(u cosx+sinx)

also given :  R = nF

so,  nu.mg = mg(u cosx+sinx)

also tanx =u

so , n = 2 (1+ u2 )-1/2

so coeff of friction: u = (1 -(4/n2) )1/2

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch